1) a) Draw a correctly labeled ASAD graph showing an economy with an output of $700,000. The output at
full potential is $800,000 and the marginal propensity to consume is 0.8.
b) Calculate the amount of government spending that would be necessary to bring the economy to long run
equilibrium. Show your work.
c) By what amount would the government need to reduce tax in order to achieve the same effect on the
economy? Explain.

d) Using the graph from part (1), show how the change in either the government spending or tax would affect the price level and output in the short-run.
Answer #1

a)

price level LRAS AS AD 700000 800000 output

b) b) spending multiplier = 1/ 1 -MPC = 1/ 1-0.8 =

spending multiplier = 1/0.2 = 5

the amount of government spending required to bring the economy to long run = 1000000/5 = 20,000

so, 20,000 government spending is required to bring the economy to long run.

c) Tax multiplier = MPC / 1 - MPC

tax multiplier = 0.8 / 0.2 = 4

government need to reduce tax in order to achieve the same effect on the economy = 1,00,000 / 4 = 25,000

d)